Jarod is having a problem with rabbits getting into his vegetable garden, so he decides to fence it in. The length of the garden is 11 feet more than 3 times the width. He needs 70 feet of fencing to do the job. What is the length?
Answer by ikleyn(52778)

.
Let W be the width in feet.  Then the length is (3W+11).


The perimeter equation is


    W + (3W+11) + W + (3W+11) = 70    feet.


    8W + 22 = 70

    8W = 70 - 22 = 48


    W = 48/8 = 6 ft     is the width.


Then the length is  3*6 + 11 = 29 ft.      Answer

Solved.